# A Unified, Practical, and Understandable Summary of Non-transactional Consistency Levels in Distributed Replication

### v1

@@ Abstract.
We present a summary of non-transactional consistency levels in the context of distributed data replication protocols. The levels are built upon a practical object pool model and are defined in a unified framework centered around the concept of ordering. We show that each consistency level can be intuitively defined by specifying two types of constraints that determine the validity of orderings allowed by the level: convergence, which bounds the lineage shape of the ordering, and relationship, which bounds the relative positions of operations in the ordering. We give examples of representative protocols and systems that implement each consistency level. Furthermore, we discuss the availability upper bound of presented consistency levels.
@@ 1. Introduction
The crucial first step towards designing distributed replication protocols and building reliable distributed storage systems is to define their consistency semantics [^1] By consistency, we refer to the constraints that restrict which orderings of operations on shared data objects are considered valid, as defined in §2. This is not to be confused with the “C” property in transactional ACID properties (Gray, 1988; Haerder and Reuter, 1983), which refers to application-level integrity invariants. In fact, consistency in our context maps to the “I” (isolation) property in ACID, as we explain in §2.. However, apart from the purely formal summary by Viotti and Vukolić (Viotti and Vukolić, 2016), there has been no unified, condensed definition of existing consistency levels in the context of distributed replication systems. This is largely due to the rich history of research that contributed to this field. Many of the fundamental works stemmed from different research areas, including distributed system modeling (Lamport, 1978b; Herlihy and Wing, 1990; Lamport, 1979; Mahajan et al., 2012; Terry et al., 1994; Lamport, 1998; Gramoli et al., 2021), multiprocessor shared-memory consistency (Hill, 1998; Mosberger, 1993; Adve and Gharachorloo, 1996; Mankin, 2007; Ahamad et al., 1995, 1993; Sorin et al., 2011), network reliability modeling (Gilbert and Lynch, 2002; Brewer, 2000; Fischer et al., 1985; Gray, 1985; Castro et al., 1999), and database transaction processing (Gray, 1988; Haerder and Reuter, 1983; Mohan et al., 1986). They discuss different pieces of the problem within different contexts, leading to plentiful but sometimes blurry terminology when applied to distributed replication.
This paper presents a unified, understandable, and sufficiently complete summary of non-transactional consistency levels in the context of a shared object store service. Compared to previous work (Viotti and Vukolić, 2016), this paper aims to achieve the following three goals. First, we propose a minimal yet self-contained theoretical framework – the Shared Object Pool (SOP) model – which unifies the definition of common consistency levels in a way that is understandable to protocol designers and system engineers. Second, we restrict our discussion to a selected set of non-transactional consistency levels seen in real object storage designs. Third, to further improve understandability, we use examples extensively to explain the practical differences between consistency levels, and refer to representative protocols and systems corresponding to each level.
Section 2 describes our problem model setup, defines logical ordering and its validity constraints, and explains the meaning of non-transactional consistency within this context. Section 3 describes all variants of ordering validity constraints. Section 4 presents the hierarchy of selected consistency levels, dissects their ordering validity constraint guarantees, explains their practical differences, and gives examples of representative protocols and systems. Section 5 discusses the availability upper bound of each level in the presence of network partitioning. Section 6 concludes this paper.
@@ 2.1. Shared Object Pool (SOP) Model
We consider a storage service shared by multiple clients, as shown in Figure 1. The service appears to be a pool of objects. Each object has a unique name and contains a value. The only way to learn about an object’s value is through the result of a client read operation, which we introduce below. Objects are not necessarily stored as physical bytes on physical machines; the SOP model is entirely conceptual and is agnostic to any actual design of protocols and implementation of systems.
Clients are single-threaded entities that invoke operations on the service. When a client $c$ issues an operation $p$, it will block until the acknowledgment of $p$ by the service. Multi-threaded or asynchronous client implementations should be modeled as multiple SOP clients.
@@ 2.2. Physical Timeline Workload
A physical timeline effectively depicts a concrete history of client activity. We can think of it as a specific “workload” that drives the storage service. Given a physical timeline, the storage service delivers a final ordering (from the set of all possible valid orderings allowed by its consistency level) that connects all operations in the timeline together.
@@ 2.3. Definition of Ordering
Given a physical timeline, an ordering is valid on the timeline with respect to a consistency level if it satisfies the validity constraints enforced by that level. We will explain validity constraints in detail in Section 3.
Early Literature Terminology. Similar definitions of “ordered before” relationship have appeared in many early literature (Lamport, 1978b, 1979; Herlihy and Wing, 1990; Attiya and Welch, 1994; Gotsman et al., 2014), where it was often termed “happens before” relationship. Unordered operations in a partial ordering were often termed “concurrent” operations. In this paper, we use “ordered before” and “happens before” interchangeably, and use “unordered” and “concurrent” interchangeably. We would like to emphasize that an ordering is a logical representation of how the storage service thinks client operations take effect. It is a completely decoupled concept from physical time.
@@ 2.4. Meaning of Consistency
An ordering represents logical dependencies among operations, similar to Lamport’s definition of logical clock (Lamport, 1978b), and does not necessarily capture physical time relationships in the timeline. In fact, whether physical time is respected or not is one of the validity constraints that differentiate several consistency levels, as we demonstrate in Section 4. Our SOP model shares similarities with the specification framework for replicated data types proposed by Burckhardt et al. (Gotsman et al., 2014); the differences are that we simplify the notion of ordering (at the cost of being less expressive in corner cases) and cover stronger consistency levels (rather than focusing only on causal and eventual consistency models).
Note that the SOP model is oblivious to any system design and implementation details of the service, including but not limited to how the service is constructed out of servers, what the network topology looks like, and how are client-server connections established. These internal design choices should not affect the interface semantic exposed to clients. Different replication protocols may make different assumptions of the system and end up providing different consistency levels and availability guarantees.
@@
Early Literature Terminology. In early literature on shared memory consistency, operations are further decomposed into events (Herlihy and Wing, 1990). The invocation and acknowledgment of an operation are considered two separate events. All events form a strictly serial sequence, named a history. Consistency levels are then defined on the validity of well-formed histories. In this paper, we simplify this notation and choose not to use the words “event” and “history”. Instead, we consider each operation $op$ as a contiguous timespan from its start (when the client issues $op$) to its end (when the service acknowledges $op$ and returns a result to the client). When discussing ordering of operations, we use partial ordering to depict incomparability if necessary, instead of merging them into a serial history of events. We found this approach easier to understand and visualize.
@@ 3.1.3. Non-convergent Partial Order (NPO)
An NPO ordering can be a partial order of operations, just like in CPO. Furthermore, reads (and RMWs) do not have to be convergent. They are allowed to only observe a subset of values from immediately-preceding operations, or apply a diverging reduction function that may produce different values on different clients given the same set of input values. Reads still have to be well-formed, meaning they cannot observe values that come from nowhere [^2] For more complex object types such as counters or queues, this means values observed must all obey return value consistency of the object semantic (Viotti and Vukolić, 2016). We assume return value consistency for all consistency levels discussed in this paper..
@@
Notice that $|e\text{R}y{:}3|$ is now allowed to only observe value 3 and know nothing about the existence of value 4; similarly for $|f\text{R}y{:}4|$.
@@ 3.2.1. Real-Time (RT)
In an RT ordering, if operation $op_{1}$ ends before operation $op_{2}$ starts in physical time (regardless of whether they come from different clients or are on different objects), then the ordering must contain $op_{1}\leadsto op_{2}$.
@@
RT is the strongest relationship constraint that any consistency level can enforce. For each client, its operations exhibit the same order as how the client issues them, because an operation naturally finishes before the start of the one that follows it on the same client. Across different clients, RT ensures that an operation observes all operations that have been acknowledged before its start.
@@ 3.2.2. Causal (CASL)
For instance, the following is an SO ordering that satisfies CASL (but does not satisfy RT), given the same example timeline presented in the RT section:
@@
The following is another valid ordering that is CPO and CASL on the same timeline example, in which case $|d\text{R}x{:}\{1,3\}|$ also observes $|e\text{W}x{\angle}3|$, setting up an additional causal dependency from $e\text{W}x{\angle}3$ to $d\text{R}x$:
@@
CASL is weaker than RT. For each client, its own operations still exhibit the same order as how the client issues them. Across different clients, however, CASL is less restrictive than RT. An operation $op_{2}$ (or a group of operations) from a client can be reordered before another operation $op_{1}$ from a different client, even though $op_{1}$ is ahead of $op_{2}$ in physical time, as long as $op_{2}$ does not causally depend on $op_{1}$. This tolerates certain forms of divergence and allows certain operations to be processed concurrently without knowing the existence of others.
@@
  • Monotonic Reads: if two reads are from the same client, then the latter read cannot observe an older state prior to what the former read has observed. This means if a client issues a read $op_{1}$ followed by another read $op_{2}$, then $op_{2}$ must be ordered after all writes that $op_{1}$ observes. In this paper, we assume a slightly stricter version of this guarantee, where $op_{2}$ must be ordered after the read $op_{1}$ itself.
  • Writes Follow Reads, i.e., Session Causality: if a client issued a read $op^{\prime}$ that observed a write $op_{1}$, and later issues a write $op_{2}$, then $op_{2}$ must become visible after $op_{1}$. In this paper, we assume a slightly stricter version of this guarantee, where $op_{2}$ must be ordered after the read $op^{\prime}$ itself [^3] Having the slightly stricter versions of Monotonic Reads and Writes Follow Reads allows us to simplify the notion of causality and use a single ordering instead of two (i.e., visibility order and arbitration order (Viotti and Vukolić, 2016)) to define the selected consistency levels..
The CASL guarantee can be defined exactly as the combination of the four session guarantees (Brzezinski et al., 2004; JEPSEN, 2016).
@@ 3.2.3. First-In-First-Out (FIFO)
For instance, the following is an SO ordering that satisfies FIFO (but does not satisfy CASL), given the same example timeline presented in the RT section:
@@
Notice that $|d\text{W}y{\angle}3|$ is now ordered before $|c\text{W}x{\angle}1|$ and $|d\text{R}x{:}1|$, breaking the causality chain. Imagine that another client $f$ is reading objects $x$ and $y$, it may then observe $d$’s write to $y$ before seeing $c$’s write to $x$. This may lead to counter-intuitive results for client applications, for example, letting a user see an updated profile page before observing that the user has been removed from the access control list.
The name FIFO comes from the fact that this level of relationship constraint is equivalent to the following definition: writes from each client must be observed by everyone in the same order as they are issued by the client. Writes from different clients are allowed to be unordered with each other. It is as if each client pushes its own writes into a separate FIFO queue.
The FIFO guarantee can be defined exactly as the combination of the Read My Writes, Monotonic Writes, and Monotonic Reads session guarantees (JEPSEN, 2016). It relaxes CASL by removing Writes Follow Reads: a write operation can now get reordered before reads that precede it on the same client, as well as any writes from other clients observed by those reads.
@@ 3.2.4. None Relationship
An ordering could of course place no restrictions on the relative positions of operations. In this case, even operations issued by the same client may get arbitrarily reordered. For example, writes by the same client may be visible to another client in a different order than issued, or that a client’s read may fail to observe its own preceding write.
@@ 4.1. Linearizability
The strongest non-transactional consistency level is linearizability, as defined by Herlihy and Wing in (Herlihy and Wing, 1990). In our model, a linearizable ordering can be defined as one that satisfies both SO and RT constraints given a physical timeline. It is a serial total order where each operation is ordered before all operations that start after its acknowledgment in real time. A service that provides linearizability is one that always gives a linearizable ordering for any physical timeline workload.
@@
State Machine Replication (SMR). Since the ordering is a serial total order, it is natural to model the object pool as a state machine and model client operations as state-transfer commands. The service acts as a coordinated set of replicated state machines (typically by replicating the log of operations) and applies committed commands in the decided serial order. This resembles the well-known State Machine Replication (SMR) approach (Schneider, 1990; Lamport, 1978a), which is widely used in modeling distributed replication systems [^4] We would like to clarify another closely related term – consensus. A consensus protocol, e.g. Paxos (Lamport, 1998, 2001) and others (Ben-Or, 1983; Aspnes, 2003), operates at a lower level than a replication protocol; it is used to achieve agreement on a single value (or a sequence of values in optimized variants) among a set of message-passing processes. An SMR protocol, e.g. Multi-Paxos (Lamport, 2001) or Raft (Ongaro and Ousterhout, 2014), typically builds on top of or inherently integrates a consensus protocol. The meaning of consensus, however, is often extended by previous literature (Ongaro and Ousterhout, 2014) to include SMR..
Our Shared Object Pool (SOP) model is equivalent to the SMR model if we put some restrictions on both sides. Specifically, an SOP model where only SO orderings are accepted is equivalent to an SMR model where the state is a collection of objects and where there are three types of commands, namely R, W, and RMW operations. Without these restrictions, on one hand, the SMR model is more expressive than the SOP model in the aspect that it allows more general state machines with custom states and custom commands, not only reads and writes. On the other hand, the SOP model is more expressive than the SMR model in the aspect that it inherently allows partial orderings, which helps us incorporate consistency levels that do not guarantee SO.

### v2

@@ Abstract.
We present a summary of practical non-transactional consistency levels in the context of distributed data replication. Unlike prior work, we build upon a simple Shared Object Pool (SOP) model and define consistency levels in a unified framework centered around the concept of ordering. This model naturally reflects modern cloud object storage services and is thus easy to understand. We show that a consistency level can be intuitively defined by specifying two types of constraints on the validity of orderings allowed by the level: convergence, which bounds the lineage shape of the ordering, and relationship, which bounds the relative positions between operations. We give examples of representative protocols and systems, and discuss their availability upper bound.
@@ 1. Introduction
A crucial step towards designing distributed replication protocols and building reliable distributed storage systems is to define their consistency semantics [^1] By consistency, we refer to the constraints that restrict which orderings of operations on shared data objects are considered valid, as defined in §2. This is not to be confused with the “C” property in transactional ACID properties (Gray, 1988; Haerder and Reuter, 1983), which refers to application-level integrity invariants. In fact, consistency in our context maps to the “I” (isolation) property in ACID, as we explain in §2.. However, apart from the purely formal summary by Viotti and Vukolić (Viotti and Vukolić, 2016), there has been no unified definition of existing consistency levels in the context of distributed replication systems. This is largely due to the rich history of research that contributed to this field. Many fundamental breakthroughs stemmed from different research areas, including distributed system modeling (Lamport, 1978b; Herlihy and Wing, 1990; Lamport, 1979; Mahajan et al., 2012; Terry et al., 1994; Lamport, 1998; Gramoli et al., 2021), multiprocessor shared-memory consistency (Hill, 1998; Mosberger, 1993; Adve and Gharachorloo, 1996; Mankin, 2007; Ahamad et al., 1995, 1993; Sorin et al., 2011), network reliability modeling (Gilbert and Lynch, 2002; Brewer, 2000; Fischer et al., 1985; Gray, 1985; Castro et al., 1999), and database transaction processing (Gray, 1988; Haerder and Reuter, 1983; Mohan et al., 1986). They discuss different pieces of the problem within different contexts, leading to plentiful but sometimes blurry terminology when applied to distributed replication.
We propose a minimal yet self-contained theoretical framework – the Shared Object Pool (SOP) model – which unifies the definition of common consistency levels in a way that is understandable to protocol designers and system engineers. We restrict our discussion to a selected set of non-transactional consistency levels seen in real object storage designs. To further improve understandability, we use examples extensively to explain the practical differences between consistency levels, and refer to representative protocols and systems corresponding to each level.
Section 2 describes our problem model setup, defines logical ordering, and explains the meaning of non-transactional consistency within this context. Section 3 defines all variants of ordering validity constraints. Section 4 presents the hierarchy of selected consistency levels, dissects their ordering validity guarantees, explains their practical differences, and gives examples of representative protocols and systems. Section 5 discusses the availability upper bound of each level in the presence of network partitioning. Section 6 concludes this paper.
@@ 2.1. Shared Object Pool (SOP) Model
We consider a storage service shared by multiple clients, as shown in Figure 1. The service appears as a pool of objects. Each object has a unique name and contains a value. The only way to learn about an object’s value is through the result of a client read operation, which we introduce below. Objects are not necessarily stored as physical bytes on physical machines; the SOP model is entirely conceptual and is agnostic to any actual design of protocols and implementation of systems.
Clients are single-threaded, closed-loop entities that invoke operations on the service. When a client $c$ issues an operation $p$, it will block until the acknowledgment of $p$ by the service. Multi-threaded or asynchronous client implementations should be modeled as multiple SOP clients.
@@ 2.2. Physical Timeline Workload
A physical timeline depicts a concrete history of client activity. We can think of it as a specific “workload” that drives the storage service. Given a physical timeline, the storage service delivers a final ordering (from the set of valid orderings allowed by its consistency level) that connects all operations in the timeline together.
@@ 2.3. Definition of Ordering
Given a physical timeline, an ordering is valid on the timeline with respect to a consistency level if it satisfies the validity constraints enforced by that level.
Early Literature Terminology. Similar definitions of “ordered before” relationship have appeared in many early literature (Lamport, 1978b, 1979; Herlihy and Wing, 1990; Attiya and Welch, 1994; Gotsman et al., 2014), where it was termed “happens before” and was associated with single-point events. Unordered events in a partial ordering were often termed “concurrent” events. In this paper, we use “ordered before” and “happens before” interchangeably, and use “unordered” and “concurrent” interchangeably, but on operations.
@@ 2.4. Meaning of Consistency
An ordering represents logical dependencies among operations, similar to Lamport’s definition of logical clock on events (Lamport, 1978b), and does not necessarily capture physical time. In fact, whether physical time is respected or not is one of the validity constraints that differentiate several consistency levels, as we demonstrate in Section 4. Our SOP model shares similarities with the specification framework for replicated data types proposed by Burckhardt et al. (Gotsman et al., 2014); the differences are that we simplify the notion of ordering and cover stronger consistency levels (rather than focusing only on causal and eventual consistency models).
Note that the SOP model is oblivious to any system design and implementation details of the service, including but not limited to how the service is constructed out of servers, what the network topology looks like, and how client-server connections are established. These internal design choices should not affect the interface semantics exposed to clients.
@@
Early Literature Terminology. In early literature on shared memory consistency, operations are further decomposed into events (Herlihy and Wing, 1990). The invocation and acknowledgment of an operation are considered two separate events. All events form a strictly serial sequence, named a history. Consistency levels are then defined on the validity of well-formed histories. In this paper, we simplify this notation and choose not to use the words “event” and “history”. Instead, we take a different approach and consider each operation $op$ as a contiguous timespan from its start (when the client issues $op$) to its end (when the service acknowledges $op$ and returns a result to the client). When discussing ordering of operations, we use partial ordering to depict incomparability if necessary, instead of merging them into a serial history of events. We found this approach easier to understand and visualize.
@@ 3.1.3. Non-convergent Partial Order (NPO)
An NPO ordering can be a partial order of operations, just like in CPO. Furthermore, reads (and RMWs) do not have to be convergent. They are allowed to only observe a subset of values from immediately-preceding operations, or apply a diverging reduction function that may produce different values on different clients given the same set of input values. Reads still have to be well-formed, meaning they cannot observe values that come from nowhere [^2] For more complex object types such as counters or queues, this means values observed must all obey return value consistency of the object semantic (Viotti and Vukolić, 2016). We assume return value consistency is held for all consistency levels discussed in this paper, as is the case in all practical cloud systems..
@@
Notice that $|e\text{R}y{:}3|$ is now allowed to only observe value 3 and miss the existence of value 4; similarly for $|f\text{R}y{:}4|$.
@@ 3.2.1. Real-Time (RT)
In an RT ordering, if operation $op_{1}$ ends before operation $op_{2}$ starts in physical time (regardless of whether they come from different clients or are on different objects), then the ordering must enforce $op_{1}\leadsto op_{2}$.
@@
RT is the strongest relationship constraint that any consistency level can enforce. For each client, its operations exhibit the same order as how the client issues them, because an operation naturally finishes before the start of the next one following it on the same client. Across different clients, RT ensures that an operation observes all other operations acknowledged before its start.
@@ 3.2.2. Causal (CASL)
For instance, the following is an SO ordering that satisfies CASL (but does not satisfy RT), given the same example timeline presented in Section 3.2.1:
@@
The following is another valid ordering that is CPO and CASL on the same timeline example; here, $|d\text{R}x{:}\{1,3\}|$ observes $|e\text{W}x{\angle}3|$, setting up an additional causal dependency from $e\text{W}x{\angle}3$ to $d\text{R}x$:
@@
CASL is weaker than RT. For each client, its own operations still exhibit the same order as how the client issues them. Across different clients, however, CASL is less restrictive. An operation $op_{2}$ (or a group of operations) from a client can be reordered before another operation $op_{1}$ from a different client, even though $op_{1}$ is ahead of $op_{2}$ in physical time, as long as $op_{2}$ has not causally observed $op_{1}$. This allows certain operations to be processed concurrently without knowing the existence of each other.
@@
  • Monotonic Reads: if two reads are from the same client, then the latter read cannot observe an older state prior to what the former read has observed. This means if a client issues a read $op_{1}$ followed by another read $op_{2}$, then $op_{2}$ must be ordered after all writes that $op_{1}$ observes.
  • Writes Follow Reads, i.e., Session Causality: if a client issued a read $op^{\prime}$ that observed a write $op_{1}$, and later issues a write $op_{2}$, then $op_{2}$ must become visible after $op_{1}$. In this paper, we assume a slightly stricter (but functionally equivalent) version of this guarantee, where $op_{2}$ must be ordered after the read $op^{\prime}$ itself [^3] Having the slightly stricter version of Writes Follow Reads allows us to simplify the notion of causality and use a single ordering instead of two (i.e., visibility order and arbitration order (Viotti and Vukolić, 2016)) to define all the selected consistency levels on the SOP model..
The CASL guarantee can be defined exactly as the combination of the above four session guarantees (Brzezinski et al., 2004; JEPSEN, 2016).
@@ 3.2.3. First-In-First-Out (FIFO)
For instance, the following is an SO ordering that satisfies FIFO (but does not satisfy CASL), given the same example timeline presented in Section 3.2.1:
@@
Notice that $|d\text{W}y{\angle}3|$ is now ordered before $|c\text{W}x{\angle}1|$ and $|d\text{R}x{:}1|$, breaking the causality chain. Imagine that another client $f$ is reading objects $x$ and $y$; it may then observe $d$’s write to $y$ before seeing $c$’s write to $x$. This may lead to counter-intuitive results for client applications, e.g., showing a user some updated private data before knowing that the user has been removed from the access control list (although the update was made after the ACL removal).
The name FIFO comes from the following analogy: writes from each client are observed by everyone in the same order as they are issued by the client, as if each client pushes its own writes into a separate FIFO queue; meanwhile, writes from different clients are not coordinated with each other by reads.
The FIFO guarantee can be defined exactly as the combination of the Read My Writes, Monotonic Writes, and Monotonic Reads session guarantees (JEPSEN, 2016). It relaxes CASL by removing Writes Follow Reads.
@@ 3.2.4. None Relationship
An ordering could of course place no restrictions on the relative positions of operations. In this case, operations issued by the same client may get arbitrarily reordered. Writes by the same client may be visible to another client in a different order than issued, and a client’s read may fail to observe its own preceding write.
@@ 4.1. Linearizability
The strongest non-transactional consistency level is linearizability, as defined by Herlihy and Wing in (Herlihy and Wing, 1990). In our model, a linearizable ordering can be defined as one that satisfies both SO and RT constraints given a physical timeline. It is a serial total order where each operation is ordered before all operations that start after its acknowledgment in real time. A service that provides linearizability is one that always yields a linearizable ordering.
@@
State Machine Replication (SMR). Since the ordering is a serial total order, it is natural to model the object pool as a state machine and model client operations as state-transfer commands. The service acts as a coordinated set of replicated state machines (typically by replicating the log of operations) and applies committed commands in the decided serial order. This resembles the well-known State Machine Replication (SMR) approach (Schneider, 1990; Lamport, 1978a), which is widely used in modeling distributed replication systems [^4] We would like to clarify another closely related term – consensus. A consensus protocol, e.g. Paxos (Lamport, 1998, 2001) and others (Ben-Or, 1983; Aspnes, 2003), operates at a lower level than a replication protocol; it is used to achieve agreement on a single value (or a sequence of values in optimized variants) among a set of message-passing processes. An SMR protocol, e.g. Multi-Paxos (Lamport, 2001) or Raft (Ongaro and Ousterhout, 2014), builds atop or inherently integrates a consensus protocol. However, previous literature often extends consensus to include SMR (Ongaro and Ousterhout, 2014)..
Our Shared Object Pool (SOP) model is equivalent to the SMR model if we put some restrictions on both sides. Specifically, an SOP model where only SO orderings are accepted is equivalent to an SMR model where the state is a collection of read-write objects. The SMR model is more expressive than the SOP model in the aspect that it allows more general state machines with custom states and custom commands, not only reads and writes. SOP is more expressive than SMR in the aspect that it inherently allows partial orderings, which helps us incorporate consistency levels that do not guarantee SO.
@@
Protocols & Systems. Linearizability is the predominant consistency level adopted by critical replication systems built atop SMR protocols. Classic protocols include Chain Replication (Renesse and Schneider, 2004), Multi-Paxos (Lamport, 2001) and its many variants/optimizations (Lamport and Massa, 2004; Lamport, 2006; Lamport et al., 2009; Howard et al., 2016; Oki and Liskov, 1988; Ongaro and Ousterhout, 2014; Mao et al., 2008; Moraru et al., 2013; Poke et al., 2017; Ailijiang et al., 2020; Enes et al., 2020; Bezerra et al., 2014; Stathakopoulou et al., 2022; Whittaker et al., 2021; Ganesan et al., 2020, 2021; Ngo et al., 2020; Mu et al., 2014; Wang et al., 2020; Fouto et al., 2022), Byzantine fault-tolerant protocols (Castro et al., 1999; Yin et al., 2019; Abd-El-Malek et al., 2005; Cowling et al., 2006), and others (Lee et al., 2015; Pan et al., 2021; Li et al., 2016; Takruri et al., 2020; Ports et al., 2015; Zhang et al., 2015) (some with advanced hardware assumptions). Systems incorporating SMR components include lock/coordination services (Burrows, 2006; Balakrishnan et al., 2013a, b), distributed cloud databases (Corbett et al., 2013; VanBenschoten et al., 2022; Huang et al., 2020; ScyllaDB, 2023; Zhou et al., 2021; Verbitski et al., 2017; Dragojević et al., 2015; rqlite, 2024), and metadata services of large-scale storage systems (Bolosky et al., 2011; Ghemawat et al., 2003; etcd, 2023; Inc., 2023; Guerraoui et al., 2022).
4.2. Sequential Consistency
Sequential consistency, as originally defined by Lamport in the context of a multiprocessor computer (Lamport, 1979), means that all clients agree on the same sequence of operations applied by the service, where operations from each client appear in the same order as issued by the client. In our model, a service that provides sequential consistency always gives an ordering that is SO and CASL [^5] Viotti and Vukolić gave a formal formula of sequential consistency that conjuncts SO with PRAM (instead of CASL as in our definition) (Viotti and Vukolić, 2016). However, we believe the formula is an erratum and deviates from their text, which reads: “the realtime ordering of operations invoked by the same process is preserved.” Their discussion indicates a conjunction with processor consistency, which aligns with our CASL constraint. for any physical timeline workload.
Compared to linearizability, since the ordering does not have to be RT, sequential consistency allows the service to move an operation (or a group of operations) backward in time, reordering it before another group that does not causally precede it. This property is sometimes referred to as unstable ordering (Bessani et al., 2010; Burke et al., 2020), in contrast to stable ordering provided by linearizability.
For example, given the following physical timeline:
$c:$$d:$$c\text{W}x{\angle}1$$d\text{W}x{\angle}2$$d\text{R}x$
A linearizable ordering must be SO and RT:
$|c\text{W}x{\angle}1|$ $|d\text{W}x{\angle}2|$ $|d\text{R}x{:}2|$
$|d\text{W}x{\angle}2|$ $|c\text{W}x{\angle}1|$ $|d\text{R}x{:}1|$
At first glance, it may be hard to tell the exact differences between linearizability and sequential consistency. Attiya and Welch presented a quantitative analysis of the performance implications of these two levels, showing that linearizability is strictly more expensive to implement than sequential consistency for common object types in systems without perfectly synchronized clocks (Attiya and Welch, 1994). But what semantic power do we lose by relaxing the real-time guarantee? The following paragraphs explain three practical implications: 1) sequential consistency does not capture external causality dependencies, 2) sequential consistency is non-local, and 3) it takes extra care to add read-modify-write (RMW) operation support to a sequentially-consistent protocol.